If the diameter of a circle is 13 x 10^-2 meters, then what is its radius?
Answer by dabanfield(803)

If the diameter of a circle is 13 x 10^-2 meters, then what is its radius?
Raduis = 1/2 od the diameter so:
Radius = (13*10^-2)/2 = (13/2)*10^-2 = 6.5*10^-2